Dr. Max ran a dentist office and he made about $120,000 a year. He had $300,000 in investments and earned a return of, on average, 12 % each year. He also earned $5,000 a year coaching the local high school basketball team.

On a whim, Dr. Max decided to change his life - he got beach fever and decided to buy and run a beach bar in Grayton Beach, Florida on the Panhandle. Of course, he had to resign as coach and give up his dental practice. He had to cash in his $300,000 to buy the bar. He is now happily selling beach drinks at his bar, and he made about $420,000 in total revenue with $270,000 in total costs (it is a popular bar. )

11) What is his accounting profit?

12) What is his economic profit?
Answer #1

Accounting profit = revenue - cost

= 420000 - 270000= 150000

Economic profit = accounting profit - opportunity cost

Opportunity cost is the income let go when he was a dentist and a coach

Opportunity cost = 120000 - 0.12(300000) + 5000= 161000

Economic profit = 150000-161000= -11000